OpenStudy (katherinesmith):

How many terms are in the expansion (x + y)^10

OpenStudy (katherinesmith):

1024?

OpenStudy (katherinesmith):

10 13 12 11 are my options

OpenStudy (raden):

sorry, my mistake. that was to determine the sum all coefficients :) now there are (n+1) terms, here n = 10 so, it is (10+1) terms = 11 terms

OpenStudy (katherinesmith):

can you help me with another?

OpenStudy (raden):

i'll try

OpenStudy (katherinesmith):

what is the 4th term of \[(a-\sqrt{2})^{8}\]

OpenStudy (raden):

it can be : 8C3 (a)^(8-4) (-sqrt(2))^3 then evaluate the calculation

OpenStudy (raden):

opsss.. i mean 8C3 (a)^(8-3) (-sqrt(2))^3

OpenStudy (raden):

you know about combination, right ?

OpenStudy (raden):

use the formula : nCr = n!/(r!(n-r)!) so, to get the exact value of 8C3, it means n=8, and r=3 8C3 = 8!/(3!(8-3)!) = 8!/(3!5!) = (8.7.6.5!)/(3.2.1.5!) cancel the 5!'s, then simplify again : = 8.7.6/3.2.1 = 56

OpenStudy (raden):

now for (-sqrt(2))^3. = (-sqrt(2)) * (-sqrt(2)) * (-sqrt(2)) = - 2 sqrt(2)

OpenStudy (raden):

therefore, the 4th term of 8C3 (a)^(8-3) (-sqrt(2))^3 is = 56 (a)^5 ( - 2 sqrt(2)) = - 112 sqrt(2) a^5